OpenStudy (anonymous):

determine whether the following are linear transformation. a)T:Mnn->R,where T(A)=tr(A) b)T:Mmn->R, where T(A)=transpose of A

OpenStudy (anonymous):

hlp pls

OpenStudy (bahrom7893):

@amistre64

OpenStudy (helder_edwin):

yes they are: let \(A=(a_{ij})\) and \(B=(b_{ij})\) (i) \[ \large \text{tr}(A+B)=\sum_{i=1}^n(A+B)_{ii}=\sum_{i= 1}^n(a_{ii}+b_{ii})\] \[ \large =\sum_{i=1}^na_{ii}+\sum_{i=1}^nb_{ii}=\text{tr}(A)+\text{tr}(B). \] and \[ \large \text{tr}(\alpha A)=\sum_{i=1}^n(\alpha A)_{ii}=\sum_{i=1}^n(\alpha a_{ii}) =\alpha\sum_{i=1}^na_{ii}=\alpha\cdot\text{tr}(A) \]

OpenStudy (helder_edwin):

quite the same for (ii)

OpenStudy (anonymous):

is it not suppose to be in the formof matrices

OpenStudy (anonymous):

@TuringTest ,hlp pls

OpenStudy (anonymous):

the 2nd one plz